OpenStudy (anonymous):

Describe the concavity of the function f(x) = 1/x^3 A. The graph is concave down for x < 0 and concave up for x > 0. B. The graph is concave up for x < 0 and concave down for x > 0. C. The graph is concave up for x < 0 and concave up for x > 0. D. The graph is concave down for x < 0 and concave down for x > 0.

OpenStudy (anonymous):

same steps as before

OpenStudy (anonymous):

The second derivative describes the concavity of a function, so first lets take it. y = f(x) = (x^2)(ln(x)) f'(x) = (2x)(ln(x)) + (x^2)(1/x) f'(x) = (2x)(ln(x)) + x f''(x) = 2ln(x) + (2x)(1/x) + 1 f''(x) = 2ln(x) + 2 + 1 f''(x) = 2ln(x) + 3 Now lets plug in 1 for x into the function. f''(1) = 2ln(1) + 3 f''(1) = 3 A concavity of POSITIVE 3 means that the graph is concave UP at x = 1. Hope that helps =)

OpenStudy (anonymous):